OpenStudy (anonymous):

Write the expression as one involving only sin (theta) and cos (theta) Sin 2(theta) + cos(theta)

OpenStudy (solomonzelman):

\[Sin^2θ+Cosθ\]

OpenStudy (solomonzelman):

What do you have to do with this?

OpenStudy (mertsj):

Use this identity: \[\sin 2\theta=2 \sin \theta \cos \theta \]

OpenStudy (solomonzelman):

hold on is it \[Sin^2θ+Cosθ~~~~~~~~or~~~~~~~~~~Sin2θ+Cosθ\]

OpenStudy (anonymous):

@SolomonZelman Sin2θ+Cosθ @mertsj how did youget that?

OpenStudy (solomonzelman):

So it's the second one?

OpenStudy (anonymous):

@SolomonZelman yes

OpenStudy (solomonzelman):

\[Use~~this:~~~~~~~~~~~~~~~~~~Sin~2θ~=~2Sin~θ~Cos~θ\]

OpenStudy (anonymous):

@solomonzelman what is that?

OpenStudy (solomonzelman):

An identity.

OpenStudy (anonymous):

i dont think my teacher taught us that identity yet

OpenStudy (solomonzelman):

So you are not allowed to use it because of that?

OpenStudy (anonymous):

oh nevermind i did learn it... well i have it written down on my plate... thanks

OpenStudy (solomonzelman):

Good.